# Any function I can actually write down is measurable, right?

### v2

@@ 4.3. What would it take to actually actually write it down?
The first issue, as already discussed, is that Jones’s construction in [11] of the $58$-unknown degree $4$ universal polynomial is not fully documented and would require a meaningful amount of work to reproduce carefully. Regardless, if our goal were to literally get a pen and a stack of paper and actually write one of these expressions down, it would be better to use the $28$-unknown degree $2\cdot 5^{60}$ polynomial shown in Figure 17, as it was chosen to optimize the number of operations and therefore literal written length of the polynomial. [^20] Or, if we were willing to go a little beyond polynomials, we could use the system of equations given in [11, Thm. 1], which uses an exponential and some binomial coefficients but is less than half as long as the system that yields Footnote 17 and only has $12$ unknowns. Clearly it would be quite easy for us to write this in place of $q$ in the polynomial given in 4.5, write four suprema and infima in front of the resulting expression, and thereby ‘actually write down’ something like the $f(x)$ in 4.8, right?
The next issue is that we would need to determine the parameters $m_{1}$, $m_{2}$, and $m_{3}$. Conceptually these code the Turing machine that enumerates the relevant computably enumerable set. Working out this Turing machine would require carefully writing out the coding of models of fragments of $\mathsf{ZFC}$ as elements of $\mathbb{N}^{\mathbb{N}}$, writing Gödel’s formula defining $L$ (which involves arithmetization of syntax, much like his proof of the incompleteness theorems), using these to define $<_{L}$, and then unwinding the various results in effective descriptive set theory we used to transfer these constructions from $\mathbb{N}^{\mathbb{N}}$ to $\mathbb{R}$ and produce the relevant semicomputable set.
@@
There is a function $g(x,y,z,n,k)$ that can be written in no more than 10 pages of standard  using integers and the functions $a+b$, $a-b$, $a\cdot b$, $\left\lfloor\frac{a}{b}\right\rfloor$, $a^{b}$, $a!$, $\max\{a,b\}$, $\min\{a,b\}$, $\gcd(a,b)$, $\mathrm{lcm}(a,b)$, and $(a\ \mathrm{mod}\ b)$ such that it is independent of $\mathsf{ZFC}$ whether $f(x)=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}\inf_{n\in\mathbb{N}}\sup_{k\in\mathbb{N}}g(x,y,z,n,k)$ is Lebesgue measurable.
@@ 5. Measurability by inspection
Definitions broadly like the one in 4.8 are fairy common in mathematics, yet it is also a relatively universal experience that, in practice, one does not need to worry about measurability of functions one has ‘actually written down.’ To what extent can we explain this phenomenon?
@@ 5.1. Measurability of analytic sets
Roughly speaking, the strongest [^22] There are of course always stronger results, such as Solovay’s result that ‘provably $\undertilde{\mathbf{\Delta}}^{1}_{2}$’ sets are Lebesgue measurable [12, Ex. 14.4]. $\mathsf{ZFC}$-provable general theorem regarding measurability is Lusin’s result that analytic sets are always universally measurable.
@@
    A function $f$ is Borel if and only if $f$ and $-f$ are both semi-$\undertilde{\mathbf{\Sigma}}^{1}_{1}$. In particular, Borel functions are semi-$\undertilde{\mathbf{\Sigma}}^{1}_{1}$.
@@
1 follows from 3 and 2.2. 2 follows immediately from 5.2. 3 follows from $f^{-1}([r,\infty])=\bigcap_{s<r,~{}s\in\mathbb{Q}}f^{-1}((r,\infty])$, $f^{-1}((r,\infty])=\bigcup_{s>r,~{}s\in\mathbb{Q}}f^{-1}([r,\infty])$ and 2.7.
@@ 5.2. An extra quantifier from $\sigma$-compactness
If $X$, $Y$, $Z$, and $W$, are Polish spaces (with $Y$ and $Z$ $\sigma$-compact) and $g:X\times Y\times Z\times W\to\overline{\mathbb{R}}$ is lower semi-continuous, then
@@ 5.3. Logical tameness and o-minimality
Another culprit in 4.8 is the interaction between quantifying over $\mathbb{R}$ and quantifying over $\mathbb{N}$. It is of course a standard fact that suprema and infima of countable families of measurable functions are measurable (since measurable sets form a $\sigma$-algebra), but it turns out that if we start from ‘logically’ tame functions, solely quantifying over $\mathbb{R}$ can be tame as well.
@@
A first-order structure $\mathcal{R}$ extending $(\mathbb{R},<)$ is o-minimal if any definable (with parameters) subset of $\mathbb{R}$ is a finite union of intervals (of possibly 00 or infinite length).
@@
Fact 5.9 (Knight, Pillay, Steinhorn [13], see also [8, Ch. 3]).
@@
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally o-minimal if there is an o-minimal expansion of the reals in which the restriction of $f$ to $[-k,k]^{n}$ is definable for each $k$ (although not necessarily uniformly).
@@
We have that $f(\bar{x})=\sup_{k\in\mathbb{N}}\sup_{\bar{y}\in[-k,k]^{m}}h(\bar{x},\bar{y})$. By local o-minimality, $\sup_{\bar{y}\in[-k,k]^{m}}h(\bar{x},\bar{y})$ is Borel for each $k\in\mathbb{N}$, whereby $f(\bar{x})$ is Borel. The infimum case follows from the supremum case by considering $-\sup_{\bar{y}\in\mathbb{R}^{m}}-h(\bar{x},\bar{y})$. ∎
Note though that in the case of real-analytic functions, 5.13 is not an improvement over 5.5.
@@ 5.4. Measurability from large cardinals
so there is some sense in which you can ‘randomly’ pick an ordinal less than $\kappa$ and almost surely find an inaccessible cardinal. I have done my best to depict a measurable cardinal in Figure 7.
Figure 7. “HIGH Up in the North in the land called Svithjod, there stands a rock. It is a hundred miles high and a hundred miles wide. Once every thousand years a little bird comes to this rock to sharpen its beak.
mmWhen the rock has thus been worn away, then a single day of eternity will have gone by.”
—Hendrik van Loon, The Story of Mankind [34]
@@
5.15 is known to not be sharp (as in weaker large cardinal hypotheses, such as the existence of a Ramsey cardinal, suffice), but measurable cardinals show up in some surprising places in mathematics. The existence of a measurable cardinal is equivalent to the existence of a discrete topological space that is not realcompact (i.e., isomorphic to a closed subspace of $\mathbb{R}^{I}$ for any $I$) and to the existence of an exact functor $F:\mathrm{Set}\to\mathrm{Set}$ that is not naturally isomorphic to the identity (as discovered independently by Trnková [32] and then later Blass [5]). It was shown by Przeździecki in [22] that the non-existence of a measurable cardinal is equivalent to the statement that every group is the fundamental group of a compact Hausdorff space. Directly concatenating this with 5.15 gives the somewhat absurd result that the existence of a group that is not the fundamental group of any compact Hausdorff space (non-vacuously) implies that the projection of the complement of the projection of any Borel subset of $\mathbb{R}^{3}$ is Lebesgue measurable. In the context of this paper more specifically, the functions in Theorems 4.8 and 4.12 are measurable if there is a measurable cardinal. A corollary of this fact (which is much easier to prove directly and was originally shown by Scott in [24]) is that there cannot be a measurable cardinal in $L$; it is ‘too thin’ to accommodate such an object. [^26] Note though that if there is a measurable cardinal $\kappa$, then the literal ordinal $\kappa$ will still exist in $L$ and will be a (fairly large inaccessible) cardinal there. It just won’t be a measurable cardinal in the sense of having a countably additive $\{0,1\}$-valued measure on its power set. There is however a modified construction, $L[U]$, which has many similar properties to $L$ and can accommodate a measurable cardinal. Silver showed in [27] that there is a $\Delta^{1}_{3}$ well-ordering of $\mathbb{R}$ in $L[U]$, [^27] See also [12, Thm. 20.18] for a modern proof of this in terms of iterable pre-mice, which are an important tool in studying the fine structure of more sophisticated inner models like $L[U]$. so we get analogs of Theorems 4.8 and 4.12 for the theory $\mathsf{ZFC}+\text{``}$ “there exists a measurable cardinal” at the cost of adding one additional supremum or infimum at the beginning of the expression defining the relevant indicator functions.
@@
By [10, Thm. 20.24, Lem. 20.25], Vopěnka’s principle implies the existence of a supercompact cardinal, which is known to imply the assumption of 2 (as discussed in [26]). The rest follows from [12, Thm. 32.9]. ∎
$L(\mathbb{R})$ is the smallest inner model containing all of the real numbers (see Fig. 8). Unlike $L$ though, it doesn’t always satisfy the axiom of choice (otherwise it would always have non-measurable sets). I would argue that $L(\mathbb{R})$ encompasses (and probably over-approximates) most mathematicians’ intuitive sense of what it means to ‘define something without the axiom of choice’ or in other words to ‘actually write something down.’ Any set of reals or real-valued function that one can prove the existence of without choice will live inside $L(\mathbb{R})$, since it is a model of $\mathsf{ZF}$. So we have that if Vopěnka’s principle holds, no analogs of Theorems 4.8 and 4.12 are possible. There can be no ‘explicit’ description of a Vitali set or a Banach–Tarski decomposition of the sphere in the presence of sufficiently strong large cardinals.
@@ 5.5. Table of (non-)measurability results
Table 1 summarizes some of the various measurability results we’ve seen in this section and compares them to the non-measurability results from Section 4 (and a couple we’ll discuss in a moment). In this table, $\operatorname*{qqq}$ represents an instance of either $\inf$ or $\sup$. The asterisks are Kleene stars, so in particular $\sup^{\ast}_{X}$ represents an arbitrarily long finite string of suprema over $X$ and $\operatorname*{qqq}^{\ast}_{X}$ represents an arbitrarily long finite string of (possibly alternating) suprema and infima. The expression $[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}$ represents an arbitrarily long finite string of $\sup_{\mathbb{R}}$’s, $\sup_{\mathbb{N}}$’s, and $\inf_{\mathbb{N}}$’s (again, possibly alternating). Of course, all of these statements entail the analogous inverted statement, with suprema and infima (and lower and upper) switched, as all of the definability classes mentioned are closed under negation (except lower semi-continuity). Finally, $\mathrm{Con}(\mathsf{ZFC})$ is the statement that $\mathsf{ZFC}$ is consistent.
@@
Of course the table is not comprehensive. For instance, by Facts 2.7 and 5.15, the presence of a measurable cardinal extends the seventh entry of the table to $[\inf_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\allowbreak[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\allowbreak\inf_{\mathbb{R},\mathbb{N}}^{\ast}\sup_{\mathbb{R},\mathbb{N}}^{\ast}g$ for continuous $g$ (with analogous extensions for the eighth and ninth entries and in the presence of stronger large cardinal assumptions). There are also more independence results as seen in the third and fourth entries. For example, it is possible (although surprisingly tedious [^29] Exercise 1. Show that for any open set $U\subseteq\mathbb{R}^{n}$, there is an entire real-analytic function $g(\bar{x},y)$ such that $\mathbf{1}_{U}(\bar{x})=\sup_{y\in\mathbb{R}}g(\bar{x},y)$ for all $\bar{x}\in\mathbb{R}^{n}$, where $\mathbf{1}_{U}(\bar{x})$ is the indicator function of $U$. For extra credit, argue informally that $g(\bar{x},y)$ can be computable if $U$ is semicomputable. Exercise 2. Show that for any open set $U_{0}\subseteq\mathbb{R}^{k}\times\mathbb{N}$, there is an open set $U\subseteq\mathbb{R}^{k+1}$ such that for any $\bar{x}\in\mathbb{R}^{k}$, $\langle\bar{x},n\rangle\in U_{0}$ for all $n\in\mathbb{N}$ if and only if $\langle\bar{x},y\rangle\in U$ for all $y\in\mathbb{R}$. Exercise 3. Use Exercises 1 and 2 to show that for any $G_{\delta}$ set $A\subseteq\mathbb{R}^{\ell}$, there is a real-analytic function $g(\bar{x},y,z)$ such that $\mathbf{1}_{A}(\bar{x})=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}g(\bar{x},y,z)$ for all $\bar{x}\in\mathbb{R}^{\ell}$. Use this and earlier results in this paper to justify the third entry of Table 1.) to produce a computable total real-analytic function $g(x,y,z,w,t)$ such that it is independent of $\mathsf{ZFC}$ whether $f(x)=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}\inf_{w\in\mathbb{R}}\sup_{t\in\mathbb{R}}g(x,y,z,w,t)$ is Lebesgue measurable, which shows that local o-minimality really doesn’t allow us to improve 5.6 in general.
@@
Function definition Definability of $g$ Measurability $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{N}}\sup_{\mathbb{N}}^{\ast}g$ Polynomial over $\mathbb{Z}$ Independent (4.8) $\sup_{\mathbb{R}}\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{N}}\sup_{\mathbb{N}}^{\ast}g$ Polynomial over $\mathbb{R}$ Implies $\mathrm{Con}(\mathsf{ZFC})$ (4.10) $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{R}}\sup_{\mathbb{R}}g$ Real-analytic Independent $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{R}}^{\ast}\sup_{\mathbb{R}}^{\ast}\inf_{\mathbb{R}}g$ Trig. polynomial Independent $\inf^{\ast}_{\mathbb{R}}\sup^{\ast}_{\mathbb{R}}\inf_{\mathbb{R}}^{\ast}\sup_{\mathbb{R}}^{\ast}g$ Trig. polynomial ????? (5.17) $[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}g$ Borel Measurable (5.4) $[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\inf_{\mathbb{R},\mathbb{N}}^{\ast}\sup_{\mathbb{R},\mathbb{N}}^{\ast}g$ Continuous Measurable (5.4, 5.6) $[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\operatorname*{qqq}_{\mathbb{R}}^{\ast}g$ O-minimal: Rational, Measurable (5.4, 5.10) Exponential, Step, etc. (e.g., $e^{x^{2}/|y+\mathbf{1}_{[3,4]}(\log z)|}$) $[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\inf_{\mathbb{R}}^{\ast}g$ Locally o-minimal: Measurable (5.4, 5.13) Real-analytic, Floor, etc. (e.g., $\lceil x^{2}\cos(e^{y\lfloor z\rfloor})\rceil$) $[\inf_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}g$ Borel Measurable cardinal (5.15) $\operatorname*{qqq}_{\mathbb{R},\mathbb{N}}^{\ast}g$ $L(\mathbb{R})$ Vopěnka’s principle (5.16)

### v3

@@ 4.3. What would it take to actually actually write it down?
The first issue, as already discussed, is that Jones’s construction in [14] of the $58$-unknown degree $4$ universal polynomial is not fully documented and would require a meaningful amount of work to reproduce carefully. Regardless, if our goal were to literally get a pen and a stack of paper and actually write one of these expressions down, it would be better to use the $28$-unknown degree $2\cdot 5^{60}$ polynomial shown in Figure 17, as it was chosen to optimize the number of operations and therefore literal written length of the polynomial. [^20] Or, if we were willing to go a little beyond polynomials, we could use the system of equations given in [14, Thm. 1], which uses an exponential and some binomial coefficients but is less than half as long as the system that yields Figure 17 and only has $12$ unknowns. Clearly it would be quite easy for us to write this in place of $q$ in the polynomial given in 4.5, write four suprema and infima in front of the resulting expression, and thereby ‘actually write down’ something like the $f(x)$ in 4.8, right?
Well, the next issue is that we would need to determine the parameters $m_{1}$, $m_{2}$, and $m_{3}$. Conceptually these code the Turing machine that enumerates the relevant computably enumerable set. Working out this Turing machine would require carefully writing out the coding of models of fragments of $\mathsf{ZFC}$ as elements of $\mathbb{N}^{\mathbb{N}}$, writing Gödel’s formula defining $L$ (which involves internal coding of syntax, much like his proof of the incompleteness theorems), using these to define $_{L}$, and then unwinding the various results in effective descriptive set theory we used to transfer these constructions from $\mathbb{N}^{\mathbb{N}}$ to $\mathbb{R}$ and produce the relevant semicomputable set.
@@
There is a function $g(x,y,z,n,k)$ that can be written in no more than 10 pages of standard  using integers and the functions $a+b$, aba-b, $a\cdot b$, $\left\lfloor\frac{a}{b}\right\rfloor$, $^{b}, a!a$!, $\max\{a,b\}$, $\min\{a,b\}$, $\gcd(a,b)$, $\mathrm{lcm}(a,b)$, and (amodb)(a\ $\mathrm${mod}\ b) such that it is independent of $\mathsf{ZFC}$ whether $f(x)=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}\inf_{n\in\mathbb{N}}\sup_{k\in\mathbb{N}}g(x,y,z,n,k)$ is Lebesgue measurable.
@@ 5. Measurability by inspection
Definitions broadly like the one in 4.8 are fairy common in mathematics, yet it is also a relatively universal experience that, in practice, one does not need to worry about measurability of functions one has ‘actually written down.’ To what extent can we explain this?
@@ 5.1. Measurability of analytic sets
Roughly speaking, the strongest [^22] There are of course always stronger results, such as Solovay’s result that ‘provably $\undertilde{\mathbf{\Delta}}^{1}_{2}$’ sets are Lebesgue measurable [15, Ex. 14.4]. $\mathsf{ZFC}$-provable general theorem regarding measurability is Lusin’s result that analytic sets are always universally measurable.
@@
    A function $f$ is Borel if and only if $f$ and f-f are both semi-$\undertilde{\mathbf{\Sigma}}^{1}_{1}$. In particular, Borel functions are semi-$\undertilde{\mathbf{\Sigma}}^{1}_{1}$.
@@
1 follows from 3 and $2.$ 2 follows immediately from 5.2. 3 follows from $f^{-1}([r,\infty])=\bigcap_{s<r,~s\in\mathbb{Q}}f^{-1}((r,\infty])$, $f^{-1}((r,\infty])=\bigcup_{s>r,~s\in\mathbb{Q}}f^{-1}([r,\infty])$ and 2.7.
@@ 5.2. An extra quantifier from $\sigma$-compactness
If $X$, $Y$, $Z$, and $W$ are Polish spaces (with $Y$ and $Z$ $\sigma$-compact) and $g:X\times Y\times Z\times W\to\overline{\mathbb{R}}$ is lower semi-continuous, then
@@ 5.3. Logical tameness and o-minimality
Another culprit in 4.8 is the interaction between quantifying over $\mathbb{R}$ and quantifying over $\mathbb{N}$. It is of course a standard fact that suprema and infima of countable families of measurable functions are measurable (since measurable sets form a $\sigma$-algebra), but it turns out that if we start from ‘logically tame’ functions, solely quantifying over $\mathbb{R}$ can be tame as well.
@@
A first-order structure $\mathcal{R}$ extending $(\mathbb{R},<)$ is o-minimal if any definable (with parameters) subset of $\mathbb{R}$ is a finite union of intervals (of possibly 0 or infinite length).
@@
Fact 5.9 (Knight, Pillay, Steinhorn [16], see also [9, Ch. 3]).
@@
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is locally o-minimal if there is an o-minimal expansion of the reals in which the restriction of $f$ to $k,k]^{n}$ is definable for each $k$ (although not necessarily uniformly).
@@
We have that $f(\bar{x})=\sup_{k\in\mathbb{N}}\sup_{\bar{y}\in[-k,k]^{m}}h(\bar{x},\bar{y})$. By local o-minimality, $\sup_{\bar{y}\in[-k,k]^{m}}h(\bar{x},\bar{y})$ is Borel for each $k\in\mathbb{N}$, whereby $f(\bar{x})$ is Borel. The infimum case follows from the supremum case by considering $\sup_{\bar{y}\in\mathbb{R}^{m}}-h(\bar{x},\bar{y})$. ∎
Note though that in the case of real-analytic functions, 5.13 is not an improvement over $5.$
@@ 5.4. Measurability from large cardinals
so there is some sense in which you can ‘randomly’ pick an ordinal less than $\kappa$ and almost surely find an inaccessible cardinal.
@@
5.15 is known to not be sharp (as in weaker large cardinal hypotheses, such as the existence of a Ramsey cardinal, suffice), but measurable cardinals show up in some surprising places in mathematics. The existence of a measurable cardinal is equivalent to the existence of a discrete topological space that is not realcompact (i.e., isomorphic to a closed subspace of $\mathbb{R}^{I}$ for any $I$) and to the existence of an exact functor $F:\mathrm{Set}\to\mathrm{Set}$ that is not naturally isomorphic to the identity (as discovered independently by Trnková [36] and then later Blass [5]). It was shown by Przeździecki in [26] that the non-existence of a measurable cardinal is equivalent to the statement that every group is the fundamental group of a compact Hausdorff space. Directly concatenating this with 5.15 gives the somewhat absurd result that the existence of a group that is not the fundamental group of any compact Hausdorff space (non-vacuously) implies that the projection of the complement of the projection of any Borel subset of $\mathbb{R}^{3}$ is Lebesgue measurable. In the context of this paper more specifically, the functions in Theorems 4.8 and 4.12 are measurable if there is a measurable cardinal. A corollary of this fact (which is much easier to prove directly and was originally shown by Scott in [28]) is that there cannot be a measurable cardinal in $L$; it is ‘too thin’ to accommodate such an object. [^26] Note though that if there is a measurable cardinal $\kappa$, then the literal ordinal $\kappa$ will still exist in $L$ and will be a (fairly large inaccessible) cardinal there. It just won’t be a measurable cardinal in the sense of having a countably additive $\{0,1\}$-valued measure on its power set. There is however a modified construction, $L[U]$, which has many similar properties to $L$ and can accommodate a measurable cardinal. Silver showed in [31] that there is a $\Delta^{1}_{3}$ well-ordering of $\mathbb{R}$ in L[U]L[U], [^27] See also [15, Thm. 20.18] for a modern proof of this in terms of iterable pre-mice, which are an important tool in studying the fine structure of more sophisticated inner models like $L[U]$. so we get analogs of Theorems 4.8 and 4.12 for the theory $\mathsf{ZFC}+\text{``}$there exists a measurable cardinal” at the cost of adding one additional supremum or infimum at the beginning of the expression defining the relevant indicator functions.
@@
By [13, Thm. 20.24, Lem. 20.25], Vopěnka’s principle implies the existence of a supercompact cardinal, which is known to imply the assumption of 2 (as discussed in [30]). The rest follows from [15, Thm. 32.9]. ∎
@@
$L(\mathbb{R})$ is the smallest inner model containing all of the real numbers (see Figure 7). Unlike $L$ though, it doesn’t always satisfy the axiom of choice (otherwise it would always have non-measurable sets). I would argue that $L(\mathbb{R})$ encompasses (and probably over-approximates) most mathematicians’ intuitive sense of what it means to ‘define something without the axiom of choice’ or in other words to ‘actually write something down.’ Any set of reals or real-valued function that one can prove the existence of without choice will live inside $L(\mathbb{R})$, since it is a model of $\mathsf{ZF}$. So we have that if Vopěnka’s principle holds, no analogs of Theorems 4.8 and 4.12 are possible. There can be no ‘explicit’ description of a Vitali set or a Banach–Tarski decomposition of the sphere in the presence of sufficiently strong large cardinals.
@@ 5.5. Table of (non-)measurability results
Table 1 summarizes some of the various measurability results we’ve seen in this section and compares them to the non-measurability results from Section 4 (and a couple we’ll discuss in a moment). In this table, $\operatorname*{qqq}$ represents an instance of either $\inf$ or $\sup$. The asterisks are Kleene stars, so in particular $\sup^{\ast}_{X}$ represents an arbitrarily long finite string of suprema over $X$ and $\operatorname*{qqq}^{\ast}_{X}$ represents an arbitrarily long finite string of (possibly alternating) suprema and infima. The expression $\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}$ represents an arbitrarily long finite string of $\sup_{\mathbb{R}}$’s, $\sup_{\mathbb{N}}$’s, and $\inf_{\mathbb{N}}$’s (again, possibly alternating). Of course, all of these statements entail the analogous inverted statement, with suprema and infima (and lower and upper) switched, as all of the definability classes mentioned are closed under negation (except lower semi-continuity). Finally, $\mathrm{Con}(\mathsf{ZFC})$ is the statement that $\mathsf{ZFC}$ is consistent.
@@
Of course the table is not comprehensive. For instance, by Facts 2.7 and 5.15, the presence of a measurable cardinal extends the seventh entry of the table to $\inf_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\allowbreak[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\allowbreak\inf_{\mathbb{R},\mathbb{N}}^{\ast}\sup_{\mathbb{R},\mathbb{N}}^{\ast}g$ for continuous $g$ (with analogous extensions for the eighth and ninth entries and in the presence of stronger large cardinal assumptions). There are also more independence results as seen in the third and fourth entries. For example, it is possible (although surprisingly tedious [^29] Exercise 1. Show that for any open set $U\subseteq\mathbb{R}^{n}$, there is an entire real-analytic function $g(\bar{x},y)$ such that $\mathbf{1}_{U}(\bar{x})=\sup_{y\in\mathbb{R}}g(\bar{x},y)$ for all $\bar{x}\in\mathbb{R}^{n}$, where $\mathbf{1}_{U}(\bar{x})$ is the indicator function of $U$. For extra credit, argue informally that $g(\bar{x},y)$ can be computable if $U$ is semicomputable. Exercise 2. Show that for any open set $U_{0}\subseteq\mathbb{R}^{k}\times\mathbb{N}$, there is an open set $U\subseteq\mathbb{R}^{k+1}$ such that for any $\bar{x}\in\mathbb{R}^{k}$, $\langle\bar{x},n\rangle\in U_{0}$ for all $n\in\mathbb{N}$ if and only if $\langle\bar{x},y\rangle\in U$ for all $y\in\mathbb{R}$. Exercise 3. Use Exercises 1 and 2 to show that for any $G_{\delta}$ set $A\subseteq\mathbb{R}^{\ell}$, there is a real-analytic function $g(\bar{x},y,z)$ such that $\mathbf{1}_{A}(\bar{x})=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}g(\bar{x},y,z)$ for all $\bar{x}\in\mathbb{R}^{\ell}$. Use this and earlier results in this paper to justify the third entry of Table 1.) to produce a computable total real-analytic function $g(x,y,z,w,t)$ such that it is independent of $\mathsf{ZFC}$ whether $f(x)=\inf_{y\in\mathbb{R}}\sup_{z\in\mathbb{R}}\inf_{w\in\mathbb{R}}\sup_{t\in\mathbb{R}}g(x,y,z,w,t)$ is Lebesgue measurable, which shows that local o-minimality really doesn’t allow us to improve 5.6 in general.
@@
Function definition Definability of $g$ Measurability $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{N}}\sup_{\mathbb{N}}^{\ast}g$ Polynomial over $\mathbb{Z}$ Independent (4.8) $\sup_{\mathbb{R}}\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{N}}\sup_{\mathbb{N}}^{\ast}g$ Polynomial over $\mathbb{R}$ Implies $\mathrm{Con}(\mathsf{ZFC})$ (4.10) $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{R}}\sup_{\mathbb{R}}g$ Real-analytic Independent $\inf_{\mathbb{R}}\sup_{\mathbb{R}}\inf_{\mathbb{R}}^{\ast}\sup_{\mathbb{R}}^{\ast}\inf_{\mathbb{R}}g$ Trig. polynomial Independent $\inf^{\ast}_{\mathbb{R}}\sup^{\ast}_{\mathbb{R}}\inf_{\mathbb{R}}^{\ast}\sup_{\mathbb{R}}^{\ast}g$ Trig. polynomial ????? (5.17) $\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}g$ Borel Measurable (5.4) $\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\inf_{\mathbb{R},\mathbb{N}}^{\ast}\sup_{\mathbb{R},\mathbb{N}}^{\ast}g$ Continuous Measurable (5.4, 5.6) $\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\operatorname*{qqq}_{\mathbb{R}}^{\ast}g$ O-minimal: Rational, Measurable (5.4, 5.10) Exponential, Step, etc. (e.g., $e^{x^{2}/|y+\mathbf{1}_{[3,4]}(\log z)|}$) $\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}\inf_{\mathbb{R}}^{\ast}g$ Locally o-minimal: Measurable (5.4, 5.13) Real-analytic, Floor, etc. (e.g., $\lceil x^{2}\cos(e^{y\lfloor z\rfloor})\rceil$) $\inf_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}[\sup_{\mathbb{R}},\operatorname*{qqq}_{\mathbb{N}}]^{\ast}g$ Borel Measurable cardinal (5.15) $\operatorname*{qqq}_{\mathbb{R},\mathbb{N}}^{\ast}g$ $L(\mathbb{R})$ Vopěnka’s principle (5.16)